# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v2

@@ 3.2 Alternating minimization algorithm
To begin, observe that with an initialized $\hat{\mathbf{w}}^{(0)}$, problem (2) is exactly the discrete optimal transport problem. Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (2) efficiently by some existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019). This corresponds to the inner optimization problem of (4). Next, we optimize (4) over $\mathbf{w}$, taking $\mathbf{\Pi}=\hat{\mathbf{\Pi}}^{(k)}$ as fixed. This procedure is repeated until convergence. Let
@@
and the procedure described above can be summarized in Algorithm 1. One advantage of the alternating minimization scheme is that the subproblems (steps (a) and (b) in Algorithm 1) are easy to optimize when $L$ satisfies some regularity conditions. Optimization guarantees are discussed in the next subsection.
@@ 4      Break if convergence criterion is met and set $\hat{\mathbf{w}}=\hat{\mathbf{w}}^{(k)}$
Next, we investigate the properties of Algorithm 1 in an important special case where closed forms can be obtained. Suppose $L(\mathbf{v}_{i},\mathbf{v}_{j})=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|^{2}$. Then it is not difficult to show that $F(\mathbf{w},\mathbf{\Pi})=\mathbf{w}^{\top}\mathbf{H}(\mathbf{\Pi})\mathbf{w}$, where $\mathbf{H}(\mathbf{\Pi})=\sum_{h=0}^{p-1}\mathbf{A}_{h}(\mathbf{\Pi})+\frac{\lambda}{2}\mathbf{I}_{n}$ and
@@
in which all unspecified entries in $\mathbf{A}_{h}(\mathbf{\Pi})$ are zero, and $\mathbf{0}_{h}$ is an $h\times h$ matrix of zeros. Hence $F$ is quadratic in $\mathbf{w}$. In fact, it can be shown that $\mathbf{H}(\mathbf{\Pi})$ is nonnegative definite for any $\mathbf{\Pi}$ satisfying the coupling constraints. Moreover, $F$ is strongly convex in $\mathbf{w}$ if $\lambda>0$ (see also Proposition 3 in Section 4). Thus many efficient solvers are applicable in this step (Boyd and Vandenberghe,, 2004; Parikh et al.,, 2014). If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e., $\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$ for some known $\mathbf{K}$ and $\mathbf{b}$, then step (b) in Algorithm 1 has the following closed form
@@
where $\mathbf{H}=\mathbf{H}(\hat{\mathbf{\Pi}}^{(k)})$.
Suppose further that $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$. That is, there is no side information about the missing values. In this case, the solution $(\hat{w}_{0},\ldots,\hat{w}_{n-1})$ to the subproblem in step (b) must satisfy the following system of linear equations which has the same number of unknowns and equations.
$\hat{w}_{s}=&\frac{\sum_{i=s\vee(p-1)}^{s+p-1}\sum_{n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}}{\frac{p\wedge(s+1)}{n_{1}-p+2}+\frac{\lambda}{2}},\quad\mbox{if }0\leq s\leq n_{1}-p+1,s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=s}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}+\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{s+p-1}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{n_{1}-s+1}{n_{1}-p+2}+\frac{s+p-n_{1}+1}{n-n_{1}-1}+\frac{\lambda}{2}},\\ &\hskip 142.26378pt\mbox{if }n_{1}-p+2\leq s\leq n_{1},s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=p-1}^{n_{1}}\sum_{j=s}^{(s+p-1)\wedge(n-1)}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{p\wedge(n-s)}{n-n_{1}-1}+\frac{\lambda}{2}},\quad\mbox{if }n_{1}+1\leq s\leq n-1,s\in\mathcal{M}\\ \hat{w}_{s}=&x_{s},\quad\mbox{if }s\notin\mathcal{M}\end{aligned}\right.$ (7)
where $\hat{\mathbf{\Pi}}=(\hat{\pi}_{ij})$ is treated as fixed. We make a few remarks regarding (7).
@@
Let us contrast TWI with the optimal interpolator. (7) shows, in general, the temporal Wasserstein imputation is a linear function of many observed data $\{x_{t}:t\notin\mathcal{M}\}$, even when $p$ is small. On the other hand, the optimal interpolator is usually a linear combination of nearby observed data when the underlying process is a finite-order AR model. For instance, suppose the underlying time series is an AR(1) process: $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $\{\epsilon_{t}\}$ is a mean-zero, unit-variance white noise. If $x_{n_{1}}=\bigstar$ is the only missing value in the time series, then the optimal interpolator (Peña and Maravall,, 1991) of $x_{n_{1}}$ is
$w_{n_{1},\mathrm{opt}}=\frac{\phi}{1+\phi^{2}}(x_{n_{1}-1}+x_{n_{1}+1}).$
In contrast, (7) shows that the temporal Wasserstein imputation (with $p=2$ and $\lambda=0$) is given by
@@
which is a convex combination of all observed values. This is somewhat expected because, as a nonparametric method, TWI uses all data to learn the underlying dynamics.
@@
(EM interpretation) Let $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ be a fixed point of the alternating minimization algorithm (Algorithm 1). That is,
$\hat{\mathbf{w}}\in\arg\min_{\mathbf{w}\in\mathcal{C}}F(\mathbf{w},\hat{\mathbf{\Pi}}),\quad\hat{\mathbf{\Pi}}\in\arg\min_{\mathbf{\Pi}}F(\hat{\mathbf{w}},\mathbf{\Pi}).$
In addition, suppose $\lambda=0$ for a moment. Let $\mathbf{u}=(u_{0},u_{1},\ldots,u_{p-1})$ be a random vector whose distribution is $\hat{\mu}_{\mathrm{pre}}$ and $\mathbf{v}=(v_{0},\ldots,v_{p-1})$ a random vector whose distribution is $\hat{\mu}_{\mathrm{post}}$. Furthermore, let the joint distribution of $(\mathbf{u},\mathbf{v})$ be given by the optimal transport coupling between $\hat{\mu}_{\mathrm{pre}}$ and $\hat{\mu}_{\mathrm{post}}$, identified as $\hat{\mathbf{\Pi}}$ (by viewing it as a joint distribution between the two discrete measures). Consider an index $s\in\mathcal{M}$ with $p-1\leq s\leq n_{1}-p+1$. By (7), it is not difficult to see that
@@
(8) shows that the temporal Wasserstein imputation $\hat{w}_{s}$ is essentially an average of conditional expectations. Thus, Algorithm 1 is closely related to the EM algorithm: To minimize the Wasserstein distance $\mathbb{E}_{\mathbf{\Pi}}\|\mathbf{u}-\mathbf{v}\|^{2}$, step (b) computes the conditional expectations under the current optimal transport coupling $\hat{\mathbf{\Pi}}$ while step (a) minimizes over the coupling space. Formula like (8) can be derived for other time indices $s$ with appropriate modifications.
Because of the optimization formulation, it is straightforward to extend TWI to multivariate time series. Suppose the multivariate time series of interest is $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$. To compute the imputation $\mathbf{W}=(\mathbf{w}_{0},\mathbf{w}_{1},\ldots,\mathbf{w}_{n-1})$ where $\mathbf{w}_{t}=(w_{t,1},\ldots,w_{t,d})^{\top}$, we may replace $\mathbf{v}_{i}(\mathbf{w})$ in (2) by $\mathbf{v}_{i}(\mathbf{W}):=(w_{i,1},\ldots,w_{i-p+1,1},w_{i,2},\ldots,w_{i-p+1,2},\ldots,w_{i-p+1,d})^{\top}$. Then Algorithm 1 can still be applied. Since the multivariate information is summarized by the distance matrix $\mathbf{L}=(l_{ij})_{0\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}$, where $l_{ij}=L(\mathbf{v}_{i}(\mathbf{W}),\mathbf{v}_{j}(\mathbf{W}))$, regardless of the dimension $d$, step (a) in Algorithm 1 is identical as the scalar case. In the notable special case where $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{2}$, (5) becomes
@@
where $\|\cdot\|_{F}$ is the Frobenius norm. Straightforward calculations show that fixing $\mathbf{\Pi}=\hat{\mathbf{\Pi}}$, if $\hat{\mathbf{W}}$ is a solution to step (b) of Algorithm 1, then for each $l\in\{1,2,\ldots,d\}$, $(\hat{w}_{0,l},\hat{w}_{1,l},\ldots,\hat{w}_{n-1,l})$ also satisfies the system of linear equations (7) in the case of no side information. Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel to lower the computational overhead and only the values of individual series are involved in computing the imputation in each component. The cross-sectional information is exploited when learning the optimal transport coupling in step (a). Thus, the algorithm is fast to implement in the multivariate setting.
We shall compare TWI with the approach proposed by Carrizosa et al., (2013), which also aims to alleviate the corrupted statistical patterns caused by imputation. Their method is based on penalizing the deviation of some moments from the theoretical values. As a result, the moments, such as the autocorrelation functions, of the imputed series are close to the pre-specified values. However, this approach has a number of drawbacks. First, good theoretical values for the moments may not be readily available in practice, especially when the missing values are many. Second, their approach promotes smoothness of the time series and hence positive autocorrelation, which does not necessarily align with the underlying time series. Finally, the optimization problem is highly non-convex, leading to an intense computational hurdle. In contrast, the proposed TWI is data-driven, completely nonparametric, and enjoys the computational advantages brought by the alternating minimization algorithm and recent advances in computational optimal transport.
@@
Next, we discuss the imputation of nonstationary time series. For an $I(1)$ process $\{x_{t}\}$, $y_{t}=x_{t}-x_{t-1}$ is stationary and one can apply imputation methods to the first-differenced series $\{y_{t}\}$. More generally, one can apply TWI to the series $y_{t}=\phi(B)x_{t}$, where $\phi(z)=(1-z)^{d}$ if $x_{t}$ is an $I(d)$ process, and $B$ is the back-shift operator. Then, an imputation for $x_{t}$ can be recovered by inverting $\phi(B)$. Note that, in this case, the number of missing data in $\{y_{t}\}$ multiplies, and we need at least $d+1$ contiguous observed values in $\{x_{t}\}$ to obtain an observed value for $\{y_{t}\}$. To mitigate the information loss, we can encode the information from the raw data via the admissible set $\mathcal{C}$. For example, consider the $I(1)$ case: If $x_{t}$ is observed, then since $y_{1}+\ldots+y_{t}=x_{t}-x_{0}$, we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed). Finally, although we have motivated TWI from the stationarity of the underlying process, TWI fares quite well with certain nonstationary processes such as cyclic time series. In Section 5, we further explore the application of TWI to nonstationary time series.

### v3

@@ 3.2 Alternating minimization algorithm
To begin, observe that with an initialized $\hat{\mathbf{w}}^{(0)}$, problem (2) is exactly the discrete optimal transport problem. Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (2) efficiently by existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019). This corresponds to the inner optimization problem of (4). Next, we optimize (4) over $\mathbf{w}$, taking $\mathbf{\Pi}=\hat{\mathbf{\Pi}}^{(k)}$ as fixed. This procedure is repeated until convergence (discussed in the next section). Let
@@
and the procedure described above can be summarized in Algorithm 1.
@@ 4   Break if convergence criterion is met and set $\hat{\mathbf{w}}=\hat{\mathbf{w}}^{(k)}$
One advantage of the alternating minimization scheme is that the subproblems (steps (a) and (b) in Algorithm 1) are easy to optimize when $L$ satisfies some regularity conditions. Suppose $L(\mathbf{v}_{i},\mathbf{v}_{j})=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|^{2}$. Then it is not difficult to show that $F(\mathbf{w},\mathbf{\Pi})=\mathbf{w}^{\top}\mathbf{H}(\mathbf{\Pi})\mathbf{w}$, where $\mathbf{H}(\mathbf{\Pi})=\sum_{h=0}^{p-1}\mathbf{A}_{h}(\mathbf{\Pi})+\frac{\lambda}{2}\mathbf{I}_{n}$ and
@@
in which all unspecified entries in $\mathbf{A}_{h}(\mathbf{\Pi})$ are zero. Hence $F$ is quadratic in $\mathbf{w}$. In fact, it can be shown that $\mathbf{H}(\mathbf{\Pi})$ is nonnegative definite for any $\mathbf{\Pi}$ satisfying the coupling constraints, and $F$ is strongly convex in $\mathbf{w}$ if $\lambda>0$ (see also Proposition 3 in Section 4). Thus many efficient solvers are applicable in this step (Boyd and Vandenberghe,, 2004). If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e., $\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$ where $\mathbf{K}$ and $\mathbf{b}$ are known, then step (b) in Algorithm 1 has the closed form
@@
where $\mathbf{H}=\mathbf{H}(\hat{\mathbf{\Pi}}^{(k)})$. Suppose further that $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$, corresponding to the case of no side information. The solution $(\hat{w}_{0},\ldots,\hat{w}_{n-1})$ to the subproblem in step (b) must satisfy the following system of linear equations.
$\left\{\begin{aligned} \hat{w}_{s}=&\frac{\sum_{i=s\vee(p-1)}^{s+p-1}\sum_{n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}}{\frac{p\wedge(s+1)}{n_{1}-p+2}+\frac{\lambda}{2}},\quad\mbox{if }0\leq s\leq n_{1}-p+1,s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=s}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}+\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{s+p-1}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{n_{1}-s+1}{n_{1}-p+2}+\frac{s+p-n_{1}+1}{n-n_{1}-1}+\frac{\lambda}{2}},\\ &\hskip 142.26378pt\mbox{if }n_{1}-p+2\leq s\leq n_{1},s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=p-1}^{n_{1}}\sum_{j=s}^{(s+p-1)\wedge(n-1)}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{p\wedge(n-s)}{n-n_{1}-1}+\frac{\lambda}{2}},\quad\mbox{if }n_{1}+1\leq s\leq n-1,s\in\mathcal{M}\\ \hat{w}_{s}=&x_{s},\quad\mbox{if }s\notin\mathcal{M}\end{aligned}\right$. (7)
where $\hat{\mathbf{\Pi}}=(\hat{\pi}_{ij})$ is treated as fixed, and for $a,b\in\mathbb{R}$, $a\wedge b=\min\{a,b\}$, $a\vee b=\max\{a,b\}$. We make a few remarks regarding (7).
@@
Let us contrast TWI with the optimal interpolator. Equation (7) shows, in general, the temporal Wasserstein imputation is a linear function of many observed data $\{x_{t}:t\notin\mathcal{M}\}$, even when $p$ is small. In contrast, the optimal interpolator is often a function of nearby observed data. For instance, suppose the underlying time series is a stationary AR(1) process: $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $\{\epsilon_{t}\}$ is a mean-zero, unit-variance white noise. If $x_{n_{1}}=\bigstar$ is the only missing value in the time series, then the optimal interpolator (Peña and Maravall,, 1991) of $x_{n_{1}}$ is $w_{n_{1},\mathrm{opt}}=\phi(x_{n_{1}-1}+x_{n_{1}+1})/(1+\phi^{2})$. In contrast, (7) shows that the temporal Wasserstein imputation (with $p=2$ and $\lambda=0$) is given by
@@
which is a convex combination of all observed values. This is expected because, as a nonparametric method, TWI uses all data to learn the underlying dynamics. Note that convex combination appears naturally in imputation problems. For instance, in synthetic control, a convex combination of untreated units is used to estimate the counterfactual (see Hur and Liang,, 2024; Abadie,, 2021).
@@
[…]
@@
Equation (8) shows that the temporal Wasserstein imputation $\hat{w}_{s}$ is essentially an average of conditional expectations. Thus, Algorithm 1 is closely related to the EM algorithm: To minimize the Wasserstein distance $\mathbb{E}_{\mathbf{\Pi}}\|\mathbf{u}-\mathbf{v}\|^{2}$, step (b) computes the conditional expectations under the current optimal transport coupling $\hat{\mathbf{\Pi}}$ while step (a) minimizes over the coupling space. Similar formulas to (8) can be derived for other time indices $s$ with appropriate modifications.
Because of the optimization formulation, it is straightforward to extend TWI to multivariate time series. Suppose the multivariate time series of interest is $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$. To compute the imputation $\mathbf{W}=(\mathbf{w}_{0},\mathbf{w}_{1},\ldots,\mathbf{w}_{n-1})$ where $\mathbf{w}_{t}=(w_{t,1},\ldots,w_{t,d})^{\top}$, we may replace $\mathbf{v}_{i}(\mathbf{w})$ in (2) by $\mathbf{v}_{i}(\mathbf{W}):=(w_{i,1},\ldots,w_{i-p+1,1},w_{i,2},\ldots,w_{i-p+1,2},\ldots,w_{i-p+1,d})^{\top}$. Then Algorithm 1 can still be applied. Since the multivariate information is summarized by the distance matrix $\mathbf{L}=(l_{ij})_{0\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}$, where $l_{ij}=L(\mathbf{v}_{i}(\mathbf{W}),\mathbf{v}_{j}(\mathbf{W}))$, regardless of the dimension $d$, step (a) in Algorithm 1 remains unchanged. In the notable special case where $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{2}$, (5) becomes
@@
where $\|\cdot\|_{F}$ is the Frobenius norm. Straightforward calculations show that fixing $\mathbf{\Pi}=\hat{\mathbf{\Pi}}$, if $\hat{\mathbf{W}}$ is a solution to step (b) of Algorithm 1, then for each $l\in\{1,2,\ldots,d\}$, $(\hat{w}_{0,l},\hat{w}_{1,l},\ldots,\hat{w}_{n-1,l})$ also satisfies the system of linear equations (7) in the case of no side information. Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel, lowering the computational overhead. Thus, the algorithm is fast to implement in the multivariate setting.
@@
Before closing this subsection, we discuss the imputation of nonstationary time series. For an $I(1)$ process $\{x_{t}\}$, $y_{t}=x_{t}-x_{t-1}$ is stationary and one can apply imputation methods to the first-differenced series $\{y_{t}\}$. More generally, one can apply TWI to the series $y_{t}=\phi(B)x_{t}$, where $\phi(z)=(1-z)^{d}$ if $x_{t}$ is an $I(d)$ process, and $B$ is the back-shift operator. Then, an imputation for $x_{t}$ can be recovered by inverting $\phi(B)$. However, in this case, the number of missing data in $\{y_{t}\}$ multiplies, and at least $d+1$ contiguous observed values in $\{x_{t}\}$ is needed to obtain an observed value for $\{y_{t}\}$. To mitigate the information loss, we can encode the information from the raw data via an admissible set $\mathcal{C}$. For example, consider the $I(1)$ case: If $x_{t}$ is observed, then since $y_{1}+\ldots+y_{t}=x_{t}-x_{0}$, we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed), which constitutes a set of linear constraints $\mathcal{C}$. Finally, although we have motivated TWI using the stationarity of the underlying process, it fares quite well with certain nonstationary processes, in addition to unit-root series, such as cyclic time series (see Section 5 for a simulation study).
3.3 Related works
We shall compare TWI with some existing nonparametric imputation approaches. Carrizosa et al., (2013) proposed a moment-matching method, which tries to minimize the discrepancies between the moments (e.g. autocorrelations) of the imputed series and some pre-specified target values. However, their non-convex algorithm is computationally intensive, and good target values for the moments may not be readily available in practice. In contrast, TWI is completely data-driven, and enjoys the computational efficiency of the alternating minimization algorithm and computational optimal transport.
Singular spectrum analysis (SSA) methods (Golyandina et al.,, 2018, 2001; Kondrashov and Ghil,, 2006) are nonparametric techniques that perform spectral analysis on the trajectory matrix of the time series $\mathbf{x}=(x_{0},x_{1},\ldots,x_{n-1})$. Specifically, the trajectory matrix is defined as $\mathbf{T}=(\mathbf{y}_{L-1},\mathbf{y}_{L},\ldots,\mathbf{y}_{n-1})$, where $\mathbf{y}_{j}=(x_{j-L+1},x_{j-L+2},\ldots,x_{j})^{\top}$ for $j=L-1,\ldots,n-1$, and $L$ is called the window length. SSA methods exploit the possible low-rank structure of $\mathbf{T}$. When $\mathbf{T}$ contains missing values, columns with missing entries are typically removed (shaped SSA, Golyandina et al.,, 2018), followed by direct or iterative imputation (Kondrashov and Ghil,, 2006). For some missing patterns this could result in too few complete columns, forcing $L$ to be small (violating a common guideline of $L\approx n/2$) and compromising the effectiveness of SSA. In contrast, TWI accommodates any missing pattern without discarding substantial data. While SSA is typically effective for imputing structural time series, TWI prevails in learning stochastic time series (see Section 5 for our simulation study). However, due to the convex nature of TWI (see Remark 1), it cannot handle series with explosive deterministic trends. Hence, it is advisable to remove the trends before applying TWI, or use SSA-based methods when the deterministic patterns are of primary interest.
A concurrent and independent work of Wang et al., (2025) proposed the Proximal Spectrum Wasserstein for Imputation (PSW-I), which, following the framework of Muzellec et al., (2020), imputes time series using optimal transport techniques. While both PSW-I and TWI focus on distribution matching, their implementations differ significantly. PSW-I uses a gradient descent algorithm that relies on automatic differentiation and some ad hoc gradient-stabilizing techniques, due to the difficulty in computing the gradient of the (regularized) Wasserstein distance with respect to the imputation. TWI, by comparison, uses a straightforward algorithm with minimal tuning. Although Wang et al., (2025) includes a convergence result, its guarantees hinge on restrictive assumptions. In contrast, TWI’s convergence and statistical properties are established under more standard conditions, as detailed in the next section.